OpenStudy (anonymous):

Are the given vectors normal? a=(5, -2) and b=(6, 15)

OpenStudy (anonymous):

It is a question on my homework.

OpenStudy (mathstudent55):

Find the two slopes. If they are negative reciprocals they are normal. 

OpenStudy (anonymous):

could you walk me through that? I'm confused

OpenStudy (mathstudent55):

Each vector has two known points. The origin (0,0) and the end point of the vector, (5, -2) for one vector and (6, 15) for the other vector.

OpenStudy (mathstudent55):

Using the two known points of each vector, find the slope of each vector. Then multiply the slopes together. If the product is -1 they are normal vectors.

OpenStudy (anonymous):

how do you find the slope?

OpenStudy (anonymous):

is it rise/run?

OpenStudy (mathstudent55):

yes difference in y over difference in x

OpenStudy (mathstudent55):

m1 = (-2 - 0)(5 - 0) m2 = (15 - 0/(6 - 0)

OpenStudy (anonymous):

oh wait I'm doing this wrong

OpenStudy (anonymous):

how do you multiply that?

OpenStudy (mathstudent55):

Sorry, I missed the fraction sign in the first slope. m1 = (-2 - 0)/(5 - 0) = -2/5 m2 = (15 - 0/(6 - 0) = 15/6 = 5/2

OpenStudy (mathstudent55):

Sorry, gtg.

OpenStudy (mathstudent55):

The slopes do multiply to -1, so they are normal.

OpenStudy (anonymous):

ok ty